Session: Feb 2006 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click

Experiment Feedback

Shanghai (Tao An at Fri Feb 17 01:44:41 2006) Failure : FORM
The formatter has a sync problem when starting EL033B. An error message "Clock is syncing" was occuring every several minutes. When the message shows, the formatter time is reseted to 2000. The power supply of formatter is normal: 5.1V and -5.0V. The problem is being checked, and we are grateful to any suggestions from VLBIers.

Torun (Kaz Borkowski at Fri Feb 17 06:41:11 2006) Minor failures : RECO
Scans between 4:00 and 5:06 lost due to problems with Mk5 recording (suspect as the cause is the power supply, that apparently couldn't feed two DPs together; presently we work with only one pack inserted).

Onsala (Michael Lindqvist at Fri Feb 17 07:47:58 2006) Success :
Cloudy weather, snow during night. Otherwise no known problems.

Urumqi (Zhang Hua at Fri Feb 17 07:56:12 2006) Minor failures :
Experiment start time at 18:35 UT,because of terminal trouble. From 23:10 to 23:46 UT Bank B is stop work,this disk label is "osod-050". I am check this disk,used "ssopen" and "SSReset" is ok.but it cann't work. So I am use a new disk. El033b used four diskes: uao-0016 osod-050 jive-029 med-0020

Medicina (C.Bortolotti at Fri Feb 17 09:06:11 2006) Success :

Westerbork (Tony Foley at Fri Feb 17 12:44:58 2006) Minor failures : PHAS
Wb used in phased array mode. tied array IF System had a problem in the 5th band with system ticks from 20:54 - 21:20UT. The system was rebooted in a gap about 21:20 and phases re-zeroed at 21:40UT. Otherwise OK.

Noto (P. Cassaro at Fri Feb 17 22:17:40 2006) Failure : FORM
From 23:21 UT to the end of the experiment the formatter has not been synchronized.

Effelsberg (dgraham at Sat Feb 18 08:07:57 2006) Minor failures : SCHED
First 2 scans lost to thunderstorm. Heavy rain particularly 22-24UT. Few scans lost for slewing

Jodrell_Bank (Alastair Gunn at Fri Mar 3 17:03:17 2006) Minor failures : ANT
The MkII telescope was used (Jb2). Telescope communication problems at about 06:00 UT. Required interface reset. No other known problems.

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)